# N-Agent Ad Hoc Teamwork

### v1

@@ A.1.1 SMAC Tasks
A.1.2 Evaluation Details
@@
For algorithm pairs that are the same, seeds are paired with themselves. For example, the seed pairs for VDN versus VDN would be $(1,1),(2,2),(3,3)$. The reason for this is that we are interested in the performance of agents when paired with known teammates. This is the model self-play score, in contrast to the algorithm self-play score (Albrecht et al., 2024).
A.1.3 Algorithm Implementation
The experiments in this paper use algorithm implementations from the ePyMARL codebase (Papoudakis et al., 2021b). The value-based methods are used without modification (i.e. IQL, VDN, QMIX), but we implement our own version of policy gradient methods (IPPO, MAPPO), based on the implementation of (Yu et al., 2022).
All methods employ recurrent actors and critics, with full parameter sharing, i.e. all agents are controlled by the same policy, where the agent id is input to the policy/critic networks, to allow behavioral differentiation between agents. For policy gradient methods, the policy architecture is two fully connected layers, followed by an RNN (GRU) layer, followed by an output layer. Each layer has 64 neurons with ReLU activation units, and employs layer normalization. The critic architecture is the same as the policy architecture. The value-based networks employ the same architecture, except that there is only a single fully connected layer before the RNN layers, and layer normalization is not used (following the ePyMARL implementation). Please consult the code base for full implementation details.
@@ A.1.3 Algorithm Implementation
Algorithm ED epochs ED Minibatches ED LR
POAM 1, 5, 10 1, 2 0.0005, 0.005
@@
A.2 Supplemental Results
A.2.1 On coordination conventions in StarCraft II
The experimental procedure detailed in Section 6.1 generates diverse teammates in SMAC by using MARL algorithms to train multiple teams of agents. Two underlying assumptions of the procedure are that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible. Both points are experimentally verified in this section.
@@ A.2.1 On coordination conventions in StarCraft II
For the SMAC tasks under consideration, teams trained using the same MARL algorithm, but different seeds, can converge to distinct coordination conventions. Figure 8 demonstrates this by depicting the return of teams trained together (matched seeds) versus those of teams that were not trained together (mismatched seeds), across all naive MARL algorithms (IPPO, IQL, MAPPO, QMIX, VDN) and tasks considered. Overall, the returns of the teams that were trained together are higher than those not trained together.
The general phenomenon has been previously observed and exploited by prior works in ad hoc teamwork (Strouse et al., 2022). This paper takes advantage of this to generate a set of diverse teammates for the StarCraft II experiments. We select tasks where the effect is significant, to ensure that there are distinct coordination behaviors for POAM to model. Tasks that were considered but subsequently ruled out for this reason include 3m vs 3m, 8m vs 8m, and 6h vs 8z.
@@
Tables 9, 10, 11, and 12 display the full cross-play results for all MARL algorithms on all tasks. Note that the tables are reflected across the diagonal axis for viewing ease. The values on the off-diagonal (i.e., where the two algorithms are not the same) are the cross-play score, while the values shown on the diagonal are self-play scores, computed as described in App. A.1.2. The cross-play and self-play scores displayed are means and standard errors. The rightmost column reflects the row average and standard deviation of the cross-play (xp) scores corresponding to the test set of VDN, QMIX, IQL, IPPO, and MAPPO, excluding the self-play score. Thus, the rightmost column reflects how well on average the row algorithm can generalize to unfamiliar teams.
@@
A.3 Cross-play performance as a function of the number of agents.
@@
A.4 Further Discussion of the Motivating Example
This section considers an extension of the motivating example discussed in Section 4; namely, whether considering longer agent modeling windows is sufficient to allow the ad hoc agents to solve the $s=1$ game without coordination.
Consider a modification of the ad hoc agents’ policies such that instead of picking the best response to teammate actions in the last step, they consider the last $t$ actions of each teammate, and pick the best response with respect to the expected actions. In this extended scenario, assume there are $M-1$ (uncontrolled) agents which always pick 00. Assume that the $M-1$ always-zero agents play without the ad hoc agent for $t$ steps. Then, the ad hoc agent $a_{h}$ is introduced into the game in the $t+1$-th step. In this scenario, $a_{h}$ will learn that all the other agents always pick zero, and will pick the optimal action $1$ accordingly.
Now consider the setting where there are $M-2$ always-zero agents playing without the ad hoc agent for $t$ steps, and two ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ are introduced into the game at the $t+1$-th step. Following the same modeling policy, the two ad hoc agents $a_{h}^{1}$ and $a_{h}^{2}$ will once again both learn that their teammates follow an always-zero policy, and therefore learn to pick the $1$ action — a suboptimal decision, as $s=2$.
In this setting, even if $a_{h}^{1}$ and $a_{h}^{2}$ modeled teammate behavior on a sliding window of length $t$ rather than every $t$ steps in batch form, they will not converge to optimal behavior. To see why, consider the actions of $a_{h}^{1}$ and $a_{h}^{2}$ in an arbitrary window of size $w$. Since they begin in unison, are deterministic, and consider the same window size, their actions will be identical for all the $t$ steps of this window. For the first $t$ steps, lacking history, both agents will pick action $1$, as we have shown above. After $k$ steps, they will both switch to 00, as we have also shown. Now, however, if they are both modeling agent action distributions using a rolling window of the same size $t$, after $t/2+1$ additional steps, they will reach a point in which both agents have chosen 00 more than 50% of the time in the last $t$ timesteps.
At that point, both agents will once again simultaneously switch their actions to $1$ to maximize the likelihood of winning the game. As a result $S$ would once again flip from 00 to $2$. Using a longer window for modeling action does not mitigate the problem of inadequately modeling other ad hoc agents, which leads to an inability to cooperate.
A.5 Computing Infrastructure
All value-based algorithms (e.g. QMIX, VDN, IQL) were run without parallelized training, while policy gradient algorithms were run with parallelized training. The servers used in our experiments ran Ubuntu 18.04 with the following configurations:

### v2

@@ A.3.2 Experimental Domains
Figure 8: Evolution of the within-episode mean squared error (left) and probability of actions that were actually taken by modelled teammates (right), computed by the POAM agent over the course of training, on the 5v6 task. The agent modeling performance improves over the course of an episode, as more data about teammate behavior is observed.
Figure 9: Learning curves of POAM and IPPO, where the value network is trained with and without uncontrolled agents’ data (UCD) on 5v6.
A.3.3 Evaluation Details
@@
For algorithm pairs that are the same, seeds are paired with themselves. For example, the seed pairs for VDN versus VDN would be $(1,1),(2,2),(3,3)$. The reason for this is that we are interested in the performance of agents when paired with known teammates. This is the model self-play score, in contrast to the algorithm self-play score [3].
A.3.4 Algorithm Implementation
The experiments in this paper use algorithm implementations from the ePyMARL codebase [29] (released under the Apache License). The value-based methods are used without modification (i.e. IQL, VDN, QMIX), but we implement our own version of policy gradient methods (IPPO, MAPPO), based on the implementation of [43].
All methods employ recurrent actors and critics, with full parameter sharing, i.e. all agents are controlled by the same policy, where the agent id is input to the policy/critic networks, to allow behavioral differentiation between agents. All methods employ the Adam optimizer to optimize all networks involved. For policy gradient methods, the policy architecture is two fully connected layers, followed by an RNN (GRU) layer, followed by an output layer. Each layer has 64 neurons with ReLU activation units, and employs layer normalization. The critic architecture is the same as the policy architecture. The value-based networks employ the same architecture, except that there is only a single fully connected layer before the RNN layers, and layer normalization is not used, following the ePyMARL implementation. Please consult the code base for full implementation details.
@@ A.3.4 Algorithm Implementation
Task Algorithm ED epochs ED Minibatches ED LR
5v6 POAM 1, 5, 10 1, 2 0.0005, 0.005
mpe-pp POAM 1, 5 1, 2 0.0005, 0.001
@@
A.4 Supplemental Results
A.4.1 Validating the Existence of Coordination Conventions
The experimental procedure detailed in Section 6.1 generates diverse teammates in SMAC and MPE by using MARL algorithms to train multiple teams of agents. Two underlying assumptions of the procedure are that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible. Both points are experimentally verified in this section.
@@ A.4.1 Validating the Existence of Coordination Conventions
For the tasks under consideration, teams trained using the same MARL algorithm, but different seeds, can converge to distinct coordination conventions. Figure 11 demonstrates this by depicting the return of teams trained together (matched seeds) versus those of teams that were not trained together (mismatched seeds), across all naive MARL algorithms (IPPO, IQL, MAPPO, QMIX, VDN) and tasks considered. Overall, the returns of the teams that were trained together are higher than those not trained together.
The general phenomenon has been previously observed and exploited by prior works in ad hoc teamwork [38]. This paper takes advantage of this to generate a set of diverse teammates for the StarCraft II experiments. We select tasks where the effect is significant, to ensure that there are distinct coordination behaviors for POAM to model. Tasks that were considered but subsequently ruled out for this reason include 3m vs 3m, 8m vs 8m, 6h vs 8z, and the MPE Spread task.
@@
Tables 12, 13, 14, 15, and 16 display the full cross-play results for all MARL algorithms on all tasks. Note that the tables are reflected across the diagonal axis for viewing ease. The values on the off-diagonal (i.e., where the two algorithms are not the same) are the cross-play score, while the values shown on the diagonal are self-play scores, computed as described in App. A.3.3. The cross-play and self-play scores displayed are means and standard errors. The rightmost column reflects the row average and standard deviation of the cross-play (xp) scores corresponding to the test set of VDN, QMIX, IQL, IPPO, and MAPPO, excluding the self-play score. Thus, the rightmost column reflects how well on average the row algorithm can generalize to unfamiliar teams.
@@
A.5 Cross-play performance as a function of the number of agents.
@@
A.6 Computing Infrastructure
All value-based algorithms (e.g. QMIX, VDN, IQL) were run without parallelizing environments, while policy gradient algorithms were run with parallelized environments. All methods were trained for 20M steps on all tasks. Each run took between 12-48 hours of compute, and used less than 2gB of GPU memory. Runs were parallelized to use computational resources most efficiently. The servers used in our experiments ran Ubuntu 18.04 with the following configurations: